OpenStudy (anonymous):

laplace transform of t^2 u(t-1). How do I make t^2 in terms of t-1?

OpenStudy (anonymous):

(t-1)^2+2(t-1)+1...?not quite sure what u r asking...

OpenStudy (anonymous):

Question asks for Laplace transform of \[f(t) = t ^{2} u(t-1)\] Hint given rewrite in terms of t-1 first

OpenStudy (anonymous):

dont think too complicatedly....just stretch it out x=t-1 f(t)=x^2u(t-1)+2xu(t-1)+u(t-1)

OpenStudy (anonymous):

f(t)=x^2u(x)+2xu(x)+u(x) sorry...

OpenStudy (anonymous):

\[\large\mathcal{L}\{t^2u_c(t-1)\}=\int_0^\infty t^2u_c(t-1)e^{-st}~dt\] Missing the subscript on the Heaviside step function, assuming that's what you're referring to.

OpenStudy (anonymous):

Ah I am still so lost in this equation :(

OpenStudy (anonymous):

Sorry, my mistake! I wasn't familiar with your notation of the Heaviside function. I know it as \(u_c(t)\), where \(c\) is the shift. Your notation is actually more obvious (a shift of 1 unit to the right), so \(u(t-1)=u_1(t)\). Sorry for the confusion.  The Laplace transform of this function is \[\mathscr{L}\{t^2u(t-1)\}=\int_0^\infty t^2u(t-1)e^{-st}~dt\] Since \(t^2u(t-1)=0\) for \(0\le t<1\) and \(t^2u(t-1)=t^2\) for \(t\ge1\), you can write the integral as \[\int_1^\infty t^2e^{-st}~dt\] This shouldn't be too difficult for you. Integrate by parts twice, or use the formula for the Laplace transform of functions of the form \(t^n\). As for the hint, I'm not really sure what "in terms of t-1" could mean.